Let a and b be integers. Prove that if a∣b, then an∣bn for all positive integers n.


1
Expert's answer
2022-04-05T06:56:00-0400

"a|b\Rightarrow \frac{b}{a}\in \mathbb{N} \Rightarrow \frac{bn}{an}=\frac{b}{a}\in \mathbb{N} \Rightarrow an|bn"